Hi,
How would I solve the following: Applications of Linear Equations

Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?
Thank you for helping me.
J.

Answer by scott8148(5891)

let d=dimes, n=nickels

"Joe has a collection of nickels and dimes that is worth $5.65" __ 10d+5n=565

"If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45"
__ 10(2d)+5(n+8)=1045 __ 20d+5n+40=1045 __ 20d+5n=1005

subtracting equations __ (20d+5n)-(10d+5n)=1005-565 __ 10d=440 __ d=44